23194 € was deposited on 4.4.2006 in a bank account. The interest rate was 3,5 % (p.a.) and the tax at source rate was 30 %. The bank added the accrued interest to the capital at the end of each year. Calculate the accumulated value of the investment on 4.4.2010 when the account was closed.


1
Expert's answer
2022-02-08T16:43:34-0500

Reasoning: Due to calculation of accumulated interest, we are therefore required to find compound interest.

Time from 4.4.2006 - 4.4.2010 = 4 years

For compound Interest, A=P(1+r÷100)n

=23194(1+3.5÷100)4

=23194×1.1475

=26615.12

I=Amount- principal

=26615.12 - 23194

I =3421.12

Tax at the source =30÷100×3421.12

=1026.34

Accrued Interest = 3421.12-1026.34

=2394.78

Accumulated Amount = 23194+2394.78

= 25588.78